One of your peerswas absent on the day of the test. The class average for 24 students was 75% before that person took the test. After the student took the test, the average increased to 76%. What did that person get on the test?
Answer by edjones(8007)

Let x be what the absent person got on the test.
(24*.75+x)/25=.76
24*.75+x=19
18+x=19
x=1.00=100%
.
Ed